If V = 2x-y+a, find E(x,y,z). Find E(1,0,1) and magnitude of E.

Grade:12

1 Answers

Samyak Jain
5 years ago
V = 2x – y + a,  here I’m asssuming a to be constant since it’s not specified.
We know E = – (\deltaV / \deltax)i –  (\deltaV / \deltay)j – (\deltaV / \deltaz)k.
\therefore  E(x,y,z)  =  – 2 i – ( – 1) j  =  – 2 i + j
i.e. here the electric field is a uniform.
So, E(1,0,1)  =  – 2 i + j
Magnitude of E is \sqrt{(-2)^2 + 1^2 +0^2} = \sqrt{5}.